# 3HDM at the ILC

### v1

@@ III.7 $A_{2}A_{2}Z$ Production : $SLD$ Mode
The $A_{2}A_{2}Z$ production channel followed by a semi-leptonic decay mode is analogous to the process discussed in Sec. III.6, yielding the same final state. Consequently, the relevant backgrounds remain unchanged and are listed in Table 13. Two benchmark mass points, denoted as $BP\text{-}11$ and $BP\text{-}12$, are considered, with their detailed properties provided in Table 16. The signal process is given by:
@@
BP-11:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma||BR$ 0.00637189 0.8188199 0.8188199 0.1 0.0002875
BP-12:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 354.62 385.08 202.86 548.99 322.35 551.64 $\sigma||BR$ 0.00661869 0.864266 0.864266 0.0673 0.0003327
Table 16: The upper portion of the table represents the mass points, BR, and total cross-section values for $BP-11$, while the lower portion represents the same for $BP-12$.
$BP-11$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 853 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 55.04 17.04 6.482
$BP-12$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 988 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 67.28 17.04 7.326
Table 17: The upper panel presents the cutflow summary for the BP-11 signal, while the lower panel shows the corresponding results for the BP-12 signal. Both cases assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.
@@ III.8 $H_{2}H_{2}^{\pm}W^{\mp}$ Production : $FHD$ Mode
Two representative benchmark points, $BP\text{-}13$ and $BP\text{-}14$, are considered in this study. Their mass configurations, BR and total production cross sections are summarized in Table 19. Given the complexity of the $4b+4j$ final state and the presence of fully hadronic decays, our selection strategy aims to keep as much signal as possible while reducing the background as much as we can. To this end, we apply the number-based selection cuts $N(b)\geq 4$ and $N(j)\geq 4$. These cuts are well-aligned with the parton-level expectations of the signal process and are effective in suppressing SM backgrounds, particularly those lacking sufficient $b$-quark content. Although the hadronic nature of the decay and the low signal cross section present challenges, we explore the discovery potential of this channel under a high-luminosity scenario. Our analysis indicates that, with an integrated luminosity of $2000~{}\text{fb}^{-1}$, the signal can be clearly discovered with a significance of $5\sigma$. The cutflow chart corresponding to this analysis, based on cut-based selections, is shown in Table 20.
Background Process Final State Cross Section($\sigma$) (pb)
@@
$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ $0.06767$
$e^{+}e^{-}\rightarrow t\bar{t}j\rightarrow b\bar{b}jjjjj$ $2b+5j$ $0.02346$
Table 18: Relevant Standard Model backgrounds for the $H_{2}H_{2}^{\pm}W^{\mp}$ (Section III.8) and $A_{2}H_{2}^{\pm}W^{\mp}$ (Section III.9) production channels in the fully hadronic decay ($FHD$) mode, featuring a final state topology of $4b+4j$.
$BP-13$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 224.16 359.96 377.98 474.54 326.93 500.67 $\sigma||BR$ 0.01095 0.8232 0.1575 1.0 0.6 0.001022
$BP-14$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 245.16 505.60 358.58 463.76 335.25 395.24 $\sigma||BR$ 0.01035 0.9159 0.1889 1.0 0.6 0.00129
Table 19: The table provides information on $H_{2}H_{2}^{\pm}W^{\mp}$ production and its subsequent decay into a $4b+4j$ final state. The upper portion of the table corresponds to $BP-13$, while the lower portion corresponds to $BP-14$.
$BP-13:$Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1533 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 104.0 302.8 5.15
$BP-14:$Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1934 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 136.4 302.8 6.51
Table 20: Cutflow charts for benchmark points $BP$-$13$ (upper table) and $BP$-$14$ (lower table). Both signals are analyzed at an integrated luminosity of $2000~{}\text{fb}^{-1}$.
@@ III.9 $A_{2}H_{2}^{\pm}W^{\mp}$ Production : $FHD$ Mode
We select two benchmark points—denoted as $BP\text{-}15$ and $BP\text{-}16$—that satisfy existing phenomenological constraints while offering representative signal kinematics. Their detailed properties, including masses, decay branching fractions, and production cross sections, are provided in Table 21. To extract the signal from the multijet-dominated environment, we impose a set of number-based selection cuts consistent with the parton-level topology: $N(b)\geq 4$ and $N(j)\geq 4$. This inclusive approach accounts for realistic detector effects, such as initial and final-state radiation and jet fragmentation, while maintaining strong discriminatory power against backgrounds lacking sufficient $b$-jet multiplicity. Despite the inherently suppressed cross sections and the challenging hadronic environment, we investigate whether this channel can yield meaningful statistical significance at a high-luminosity collider setup. Our results, summarized in the cutflow Table 22, indicate that with an integrated luminosity of $2000~{}\text{fb}^{-1}$, the signal can be probed with a discovery-level sensitivity of $5\sigma$.
BP-15:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 365.69 370.25 222.74 366.90 309.51 437.50 $\sigma||BR$ 0.01111 0.96079 0.28799 1.0 0.6 0.002214
BP-16:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 286.48 386.58 204.09 360.96 300.06 401.01 $\sigma||BR$ 0.011191 0.95317 0.24065 1.0 0.6 0.001848
Table 21: The upper table presents the mass points, branching ratios ($BR$), and total cross section for $BP-15$, while the lower table provides the corresponding information for $BP-16$.
$BP-15:$ Cuts Signal (S) Background (B) S vs B
Initial (no cut) 3320 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 110.24 302.8 5.43
$BP-16:$ Cuts Signal (S) Background (B) S vs B
Initial (no cut) 2772 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 143.73 302.8 6.8
Table 22: The upper portion of the table presents the cutflow chart for $BP\text{-}15$, while the lower portion corresponds to $BP\text{-}16$. At an integrated luminosity of $2000~{}\text{fb}^{-1}$.
@@ IV Discussion and Conclusions
We performed a detailed analysis of several production channels at the ILC, including $e^{+}e^{-}\to H_{2}A_{2}$, $e^{+}e^{-}\to H_{2}H_{2}Z$, $e^{+}e^{-}\to A_{2}A_{2}Z$, $e^{+}e^{-}\to H_{2}H_{2}^{\pm}W^{\mp}$, $e^{+}e^{-}\to A_{2}H_{2}^{\pm}W^{\mp}$, and $e^{+}e^{-}\to H_{1}H_{2}A_{2}$, which yield final states containing multiple jets, leptons, and b-quarks, making them ideal channels to investigate the extended Higgs sector. For each channel, we selected representative benchmark points and conducted detailed phenomenological analysis. It was observed that, with the exception of the $e^{+}e^{-}\to H_{2}A_{2}$ channel, the application of additional kinematic variables such as $p_{T}(j_{1})$ or $p_{T}(b_{1})$ beyond the required object multiplicity cuts offers limited improvement in signal sensitivity. This is due to significant overlap between signal and background distributions after basic selection, making further discrimination through such observables less effective.
This study presents a phenomenological survey of several channels through which new physics interaction vertices can be probed. In all the processes considered, at least two new particles contribute to the final state. For example, in $e^{+}e^{-}\to H_{2}A_{2}$, both $H_{2}$ and $A_{2}$ decay into $b$-quark pairs with overlapping kinematics due to their similar masses, making it difficult to reconstruct them individually from $b$-jet pairings. Nonetheless, an excess of events can still reveal the presence of such interactions. Similar challenges arise in all other channels. While addressing these combinatorial issues is beyond the scope of this work, future studies may explore dedicated strategies for improving signal reconstruction in such scenarios [25, 26]. However, the channels considered in our study do exhibit promising discovery potential at various luminosity thresholds. With the application of suitable number-based selection cuts and transverse momentum requirements on the leading and subleading $b$-quarks, the $H_{2}A_{2}$ channel yields a $5\sigma$ discovery significance at an integrated luminosity of $100~{}\text{fb}^{-1}$ for both benchmark points BP-1 and BP-2. For the $h_{1}H_{2}A_{2}$ channel, discovery is achievable at $1000~{}\text{fb}^{-1}$ for BP-3 and at $1200~{}\text{fb}^{-1}$ for BP-4. Similarly, the $H_{2}H_{2}Z$/$A_{2}A_{2}Z$ channel in the FHD mode requires approximately $300~{}\text{fb}^{-1}$ for discovery, whereas in the SLD mode, it requires about $2000~{}\text{fb}^{-1}$. Finally, the channel involving a charged Higgs, $H_{2}H_{2}^{\pm}W^{\mp}$/$A_{2}H_{2}^{\pm}W^{\mp}$, in the FHD mode demands a luminosity of up to $2000~{}\text{fb}^{-1}$ to reach discovery sensitivity. These results demonstrate that several production channels hinting at new physics can achieve a $5\sigma$ discovery at this collider setup.
@@ V Acknowledgements
A.K. acknowledges the support from the Director’s Fellowship at IIT Gandhinagar. G.B.K. acknowledges the initial financial support received under the Prime Minister’s Research Fellowship (Grant No. PMRF-192002-1802), awarded by the Ministry of Education, Government of India.

### v2

@@ III.7 $A_{2}A_{2}Z$ Production : $SLD$ Mode
The $A_{2}A_{2}Z$ production channel followed by a semi-leptonic decay mode is analogous to the process discussed in subsection III.6, yielding the same final state. Consequently, the relevant backgrounds remain unchanged and are listed in Table 13. A benchmark mass point, denoted as BP6, is considered, with its detailed properties provided in Table 16. The signal process is given by:
@@
BP6:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma\times BR$ 0.00637189 0.8188199 0.8188199 0.0673 0.0002875
Table 16: The table represents the mass points, BR, and total cross-section values for BP6.
BP5 : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 575 22147
SEL: $N(b)\geq 4$ &\& $N(\ell)\geq 2$ 37.07 17.04 5.04
Table 17: The table presents the cutflow summary for the BP6 signal. Here we assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.
@@ III.8 $H_{2}H_{2}^{\pm}W^{\mp}$ Production : $FHD$ Mode
A representative benchmark point, BP7 is considered in this study. Their mass configurations, BR and total production cross sections are summarized in Table 19. Given the complexity of the $4b+4j$ final state and the presence of fully hadronic decays, our selection strategy aims to keep as much signal as possible while reducing the background as much as we can. To this end, we apply the number-based selection cuts $N(b)\geq 4$ and $N(j)\geq 4$. These cuts are well-aligned with the parton-level expectations of the signal process and are effective in suppressing SM backgrounds, particularly those lacking sufficient $b$-quark content. Although the hadronic nature of the decay and the low signal cross section present challenges, we explore the discovery potential of this channel under a high-luminosity scenario. Our analysis indicates that, with an integrated luminosity of $2000~\text{fb}^{-1}$, the signal can be clearly discovered with a significance of $5\sigma$. The cutflow chart corresponding to this analysis, based on cut-based selections, is shown in Table 21.
Background Process Final State Cross Section $\sigma$ (pb)
@@
$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ $6.767\times 10^{-2}$
$e^{+}e^{-}\rightarrow t\bar{t}j\rightarrow b\bar{b}jjjjj$ $2b+5j$ $2.346\times 10^{-2}$
$e^{+}e^{-}\rightarrow t\bar{t}jj\rightarrow b\bar{b}jjjjjj$ $2b+6j$ $5.821\times 10^{-3}$
$e^{+}e^{-}\rightarrow t\bar{t}jjj\rightarrow b\bar{b}jjjjjjj$ $2b+7j$ $1.098\times 10^{-3}$
Table 18: Relevant Standard Model backgrounds for the $H_{2}H_{2}^{\pm}W^{\mp}$ (subsection III.8) and $A_{2}H_{2}^{\pm}W^{\mp}$ (subsection III.9) production channels in the fully hadronic decay (FHDFHD) mode, featuring a final state topology of $4b+4j$.
BP7:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 224.16 359.96 377.98 474.54 326.93 500.67 $\sigma\times BR$ 0.01095 0.8232 0.1575 1.0 0.6 0.001022
Table 19: The table provides information on $H_{2}H_{2}^{\pm}W^{\mp}$ production and its subsequent decay into a $4b+4j$ final state for BP7.
@@ III.9 $A_{2}H_{2}^{\pm}W^{\mp}$ Production : $FHD$ Mode
We select a benchmark point—denoted as BP8—that satisfy existing phenomenological constraints while offering representative signal kinematics. Their detailed properties, including masses, decay branching fractions, and production cross sections, are provided in Table 20. To extract the signal from the multijet-dominated environment, we impose a set of number-based selection cuts consistent with the parton-level topology: $N(b)\geq 4$ and $N(j)\geq 4$. This inclusive approach accounts for realistic detector effects, such as initial and final-state radiation and jet fragmentation, while maintaining strong discriminatory power against backgrounds lacking sufficient $b$-jet multiplicity. Despite the inherently suppressed cross sections and the challenging hadronic environment, we investigate whether this channel can yield meaningful statistical significance at a high-luminosity collider setup. Our results, summarized in the cutflow Table 22, indicate that with an integrated luminosity of $2000~\text{fb}^{-1}$, the signal can be probed with a discovery-level sensitivity of $5\sigma$.
BP8:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 365.69 370.25 222.74 366.90 309.51 437.50 $\sigma\times BR$ 0.01111 0.96079 0.28799 1.0 0.6 0.002214
Table 20: The table represents the mass points, branching ratios, and total cross-section values for BP8 for the $A_{2}H_{2}^{\pm}W^{\mp}\rightarrow 4b+4j$ final state.
BP7 : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 2043 198997
SEL: $N(b)\geq 4$ &\& $N(j)\geq 4$ 104.0 302.8 5.15
Table 21: Cut-flow results for benchmark points BP7. The signal and background events are analyzed at an integrated luminosity of $2000~\text{fb}^{-1}$.
BP8 : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 4428 198997
SEL: $N(b)\geq 4$ &\& $N(j)\geq 4$ 110.24 302.8 5.43
Table 22: Cut-flow results for benchmark points BP8 at an integrated luminosity of $2000~\text{fb}^{-1}$.
@@ IV Discussion and Conclusions
We performed a detailed analysis of several production channels at the ILC, including $^{+}e^{-}\to H_{2}A_{2}, e+e$$^{+}e^{-}\to H_{2}H_{2}Z$, $^{+}e^{-}\to A_{2}A_{2}Z, e+e$$^{+}e^{-}\to H_{2}H_{2}^{\pm}W^{\mp}$, $e^{+}e^{-}\to A_{2}H_{2}^{\pm}W^{\mp}$, and $e^{+}e^{-}\to H_{1}H_{2}A_{2}$, which yield final states containing multiple jets, leptons, and b-quarks, making them ideal channels to investigate the extended Higgs sector. For each channel, we selected representative benchmark points and conducted detailed phenomenological analysis. It was observed that, with the exception of the $e^{+}e^{-}\to H_{2}A_{2}$ channel, the application of additional kinematic variables such as $p_{T}(j_{1})$ or $p_{T}(b_{1})$ beyond the required object multiplicity cuts offers limited improvement in signal sensitivity. This is due to significant overlap between signal and background distributions after basic selection, making further discrimination through such observables less effective.
This study presents a phenomenological survey of several channels through which new physics interaction vertices can be probed. In all the processes considered, at least two new particles contribute to the final state. For example, in $e^{+}e^{-}\to H_{2}A_{2}$, both $H_{2}$ and $A_{2}$ decay into $b$-quark pairs with overlapping kinematics due to their similar masses, making it difficult to reconstruct them individually from $b$-jet pairings. Nonetheless, an excess of events can still reveal the presence of such interactions. Similar challenges arise in all other channels. While addressing these combinatorial issues is beyond the scope of this work, future studies may explore dedicated strategies for improving signal reconstruction in such scenarios [40, 41]. However, the channels considered in our study do exhibit promising discovery potential at various luminosity thresholds. With the application of optimized selection cuts and transverse momentum requirements on the leading and subleading $b$-jets, the $H_{2}A_{2}$ channel achieves a $5\sigma$ discovery significance at an integrated luminosity of approximately $100~\text{fb}^{-1}$ for benchmark points BP1. For the $h_{1}H_{2}A_{2}$ channel, discovery is attainable at $1000~\text{fb}^{-1}$ for BP2. In the $H_{2}H_{2}A_{2}$ channel, the benchmark point BP3 require $500~\text{fb}^{-1}$, while BP5 need around $3000~\text{fb}^{-1}$ for discovery. The $A_{2}H_{2}A_{2}$ channel reaches $5\sigma$ at $400~\text{fb}^{-1}$ for BP4, and $2000~\text{fb}^{-1}$ for BP6. Finally, channels involving a charged Higgs boson, $H_{2}H_{2}^{\pm}W^{\mp}$ and $A_{2}H_{2}^{\pm}W^{\mp}$, demand an integrated luminosity close to $2000~\text{fb}^{-1}$ to achieve $5\sigma$ discovery significance for BP7 and BP8. These results collectively demonstrate that multiple new physics channels within this framework can reach discovery sensitivity at the considered collider setup.
@@ V Acknowledgements
A.K. acknowledges the support from the Director’s Fellowship at IIT Gandhinagar. G.B.K. acknowledges the initial financial support received under the Prime Minister’s Research Fellowship (Grant No. PMRF-192002-1802), awarded by the Ministry of Education, Government of India. We thank the anonymous referee for a careful reading of the manuscript and for several helpful comments and suggestions.